Physics from CDF II in 2000-5

Pekka Sinervo University of Toronto

• The CDF II Experiment

• IPP role in CDF II

• Physics Prospects for CDF II • Schedule for Run II

IPP Planning Meeting 3 June 2000

The CDF II Experiment

CDF II gives us access to the highest-energy frontier

• Provide pp̄ collisions at √s = 2.0 TeV

• Luminosities of L ∼ 2 × 10^{32} cm^{−2}s^{−1}

• Total integrated luminosity of 5 − 10 fb^{−1}

Canadian scientists have played a strong role in CDF instrumentation, data analysis and physics

• Run I contributions – DAQ systems

– Offline software and reconstruction – Top and bottom quark physics

• Poised to continue pursuing this science with CDF II

• Represents a 4-5 year commitment

Run I Physics Contributions

IPP group had a focussed science agenda for Run I

• Top quark physics – Top quark discovery and mass

– Single top production (talk next week!)

– Top quark Pt distribution (talk next week!)

• B quark production and fragmentation – b → B^{+} , B^{0}, B_{s}, Λ_{b} rates

– B → ψ(2S) decays

– Charmonium polarization (talk next week!)

• High P_{T} Physics – Search for W′ → μν_{μ}

– Study M(jj) in high mass dijet data

Top Quark Mass

Measurements now limited by statistical uncertainties

• Overall systematic uncertainty is ∼ 5 GeV/c^{2}

– Dominated by uncertainties in soft gluon contributions

and energy scale

• Run I statistics limit systematic uncertainty

The final Run I values are

M_{top} = 176.0 ± 6.5 GeV c^{2} (CDF)
M_{top} = 172.1 ± 7.1 GeV c^{2} (D0 / )
M_{top} = 174.3 ± 5.1 GeV c^{2} (RunI)

Latest W boson mass results

M_{W} = 80.433 ± 0.079 GeV c^{2} CDF
M_{W} = 80.474 ± 0.093 GeV c^{2} D0
M_{W} = 80.350 ± 0.056 GeV c^{2} LEP2

Top Quark Transverse Momentum

Measured P_{T} smearing using MC/detector simulation

• Use standard HERWIG MC calculations

• Most smearing arises from incorrect jet assignments

Top Quark Differential Cross Section

R_{1} + R_{2} = 0.66_{−0.17}^{+0.17} (stat) _{−0.06}^{+0.06} (syst)
R_{4} < 0.16 at 95% C.L.

P_{T} Bin  Measured Fraction of Top Quarks
0 < P_{T} < 75 GeV/c  R_{1} = 0.21_{−0.21}^{+0.22} (stat) _{−0.08}^{+0.10} (syst)
75 < P_{T} < 150 GeV/c  R_{2} = 0.45_{−0.23}^{+0.23} (stat) _{−0.07}^{+0.04} (syst)
150 < P_{T} < 225 GeV/c  R_{3} = 0.34_{−0.12}^{+0.14} (stat) _{−0.05}^{+0.07} (syst)
225 < P_{T} < 300 GeV/c  R_{4} = 0.000_{−0.000}^{+0.031} (stat) _{−0.000}^{+0.019} (syst)

Run I Record

9 Ph.D. students will have completed studies

• G. Sganos on B branching fractions

• A. Warburton on B → ψ(2S)X rates

• W. Taylor on B fragmentation using semileptonic decays • H. Kim on W′ → μν_{μ} search

• K. Kordas on B Penguin decays

• B. Hinrichsen on high mass dijet events

• R. Cropp on J/ψ polarization

• A. Robinson on Top quark P_{T} distributions

• M. Popovic on B_{s} lifetimes (in progress)

Primary authors for 2 PRD’s, 3 PRL’s and 2 more in pipeline...

• Major contributions to many others

Overall scientific impact is high

Run II Forecasts

Run II will start this summer at √s = 2.0 TeV

• CDF II will begin with commissioning run (no SVX II)

• Will be completed and rolled-in by March 2001

• Physics running for first two years (IIa) ⇒ 2 fb^{−1}

• Possible goal of 15 fb^{−1} by 2006 (IIb)

Will be worth the wait

• Begin with ∼ 1 fb^{−1}/year per detector

• Finish up with ∼ 5 fb^{−1}/year in 2005-6

• Literal “Top Factory” – in 2 years, – 300 dilepton events

– 3000 tagged lepton+jet events – 1500 events in “mass sample”

IPP Physics Goals for Run II

Mass analysis uncertainties scale with statistics

• Can expect δM_{t} ∼ 3.0˜ GeV/c^{2}

• With δM_{W} ∼ 40 MeV/c^{2}

\frac{δM_{Higgs}}{M_{Higgs}} ∼ 40%

• Powerful “pointer” for LHC searches

Single top production

• Measure σ_{t} with 10% uncertainty

• Measure V_{tb} with 12% uncertainty

CP violation studies in the B sector

• Very competitive in B^{0} sector with BaBar

• Only way of directly measuring in B_{s} sector – will be able to cover up to x_{s} < ∼ 60

Planned IPP Involvement

Effective involvement in CDF II programme for 3 years

• Responsible for Si detector positioning system – Designed “inchworm” system and controls – Commitment to SVX and ISL alignment

• Lead role in offline reconstruction and simulation – Responsible for coordination of all calorimetry – Are developing

∗ GEANT/GFLASH simulation

∗ Calorimetry and jet reconstruction

Expect to continue these responsibilities

• Currently are commissioning inchworms

• Preparing for “Mock Data Challenge 2”

Personnel and Resources

Involvement expected to remain constant for Run IIa

• Currently Trischuk, Sinervo and Orr – Scope for additional involvement

• Three research associates

• 5-6 graduate students (may increase)

Expect some evolution in Run IIb

• Almost totally operational and data analysis phase

• But will remain primary training ground for next generation of experimentalists

Funding requirements will not grow significantly

• Requires $250K/annum operating support over next 4 years • Expect to ramp down after 2005

Summary

CDF-II will be one of the most exciting particle physics project in the next 5 year period

• IPP has significant investment in exploring high-energy frontier

• IPP is well-positioned to continue lead role

• Strong team of young (relatively!) scientists

Strategically, provides IPP community with early window into LHC physics

• Top and bottom quark sectors are clear focus

• Will face technological barriers similar to ATLAS

• Fills in the “physics gap” that overwise is created by lengthy lead-in to LHC programme